# SN 2024aecx: A double-peaked rapidly evolving Type IIb supernova at 11 Mpc

### v1

@@ Acknowledgments
Mephisto is developed at and operated by the South-Western Institute for Astronomy Research of Yunnan University (SWIFAR-YNU), funded by the “Yunnan University Development Plan for World-Class University” and “Yunnan University Development Plan for World-Class Astronomy Discipline”. The authors acknowledge support from the “Science & Technology Champion Project” (202005AB160002) and from two “Team Projects” – the “Top Team” (202305AT350002) and the “Innovation Team” (202105AE160021), all funded by the “Yunnan Revitalization Talent Support Program”. We also acknowledge the National Key Research and Development Program of China (2024YFA1611603). BK is supported by the “Special Project for High-End Foreign Experts”, Xingdian Funding from Yunnan Province. JZ is supported by the National Natural Science Foundation of China (12173082, 12333008), the Yunnan Fundamental Research Projects (grants 202501AV070012, 202401BC070007 and 202201AT070069), the Top notch Young Talents Program of Yunnan Province, the Light of West China Program provided by the Chinese Academy of Sciences, the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001). We acknowledge the support of the staff of the LJT. Funding for the LJT has been provided by the CAS and the People’s Government of Yunnan Province. The LJT is jointly operated and administrated by YNAO and the Center for Astronomical Mega-Science, CAS. We also acknowledge WISeREP - https://www.wiserep.org.
This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile.

### v2

@@ Acknowledgments
Mephisto is developed at and operated by the South-Western Institute for Astronomy Research of Yunnan University (SWIFAR-YNU), funded by the “Yunnan University Development Plan for World-Class University” and “Yunnan University Development Plan for World-Class Astronomy Discipline”. The authors acknowledge support from the “Science & Technology Champion Project” (202005AB160002) and from two “Team Projects” – the “Top Team” (202305AT350002) and the “Innovation Team” (202105AE160021), all funded by the “Yunnan Revitalization Talent Support Program”. We also acknowledge the National Key Research and Development Program of China (2024YFA1611603). BK is supported by the “Special Project for High-End Foreign Experts”, Xingdian Funding from Yunnan Province. JZ is supported by the National Natural Science Foundation of China (12173082, 12333008), the Yunnan Fundamental Research Projects (grants 202501AV070012, 202401BC070007 and 202201AT070069), the Top notch Young Talents Program of Yunnan Province, the Light of West China Program provided by the Chinese Academy of Sciences, the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001).
We thank the staff of IAO, Hanle, and CREST, Hosakote, that made HCT observations possible. The facilities at IAO and CREST are operated by the Indian Institute of Astrophysics, Bangalore. We acknowledge the support of the staff of the LJT. Funding for the LJT has been provided by the CAS and the People’s Government of Yunnan Province. The LJT is jointly operated and administrated by YNAO and the Center for Astronomical Mega-Science, CAS. We also acknowledge WISeREP - https://www.wiserep.org.
This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. We also thank the referee for insightful comments and suggestions that significantly improved the manuscript.